OpenStudy (decarr432):

Given that 12,3 is on the graph of f(x), Find the corresponding point of the function 8f(x)

OpenStudy (danjs):

notice you arent changing the x value here, it is still just f(x), but you are multiplying the y value by that constant 8 points (x , f(x))

OpenStudy (decarr432):

So would it be 12,24?

OpenStudy (danjs):

yes, any x value you pick will still be evaluated in y = f(x), except the value is now y = 8*f(x)

OpenStudy (decarr432):

Okay thank you very much Iv'e been trying to take this test on plato so i'm going over the tutorial and pretest

OpenStudy (danjs):

you can test it with a made up thing like y = f(x) = x^2 and y = 8*f(x) = 8*x^2

OpenStudy (danjs):

welcome